OpenStudy (anonymous):

What is the solution of the equation? 3x²=100 a: -9\25, 9\25 b: -3\5, 3\5 c: -25\9, 25\9 d: -5\3, 5\3

OpenStudy (anonymous):

x = 10/sqrt(3)

OpenStudy (ddcamp):

3x² = 100 x² = (100/3) \[x = \pm\sqrt{\frac{ 100 }{ 3 }} = \pm\frac{ \sqrt{100} }{ \sqrt{3} } = \pm\frac{ 10 }{ \sqrt{3} } = \pm\frac{ 10\sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

@DDCamp how would I do that with no graphing calculator?

OpenStudy (anonymous):

graphing?

OpenStudy (anonymous):

you just need mental math

OpenStudy (anonymous):

well I don't know the sqrt of 33.33